OpenStudy (anonymous):

To solve the problem BELOW, find the angle in the interval [ -π/2 , π/2 ] whose tangent equals 0.

OpenStudy (anonymous):

\[\tan^{-1}(\cos \frac{ π }{ 2 })\]

OpenStudy (anonymous):

I know that cos(π/2) = 0º

OpenStudy (psymon):

So what angle does tan = 0?

OpenStudy (anonymous):

Does that mean:\[\tan^{-1}(0) = 0\]

OpenStudy (anonymous):

You said what angle...\[\tan(0º) = 0\]

OpenStudy (psymon):

Right, 0, lol.

OpenStudy (anonymous):

Wow, thats it? I thought there would be more to the problem...

OpenStudy (anonymous):

@radar @genius12 @amistre64 @phi Psymon is offline now, and I kinda fell asleep before I asked him, but is that all to my problem?

OpenStudy (phi):

cos(pi/2) = 0 tan(0) is 0 you should know the sin,cos,tan of special angles 0,30,45,60 and 90 degrees

OpenStudy (phi):

knowing tan(0)= 0 you also know \[ \tan^{-1}(\tan(0))= \tan^{-1}(0) \\ 0 = \tan^{-1}(0) \\ \tan^{-1}(0)=0\]

OpenStudy (anonymous):

Gotcha!

OpenStudy (amistre64):

cos pi/2 is a ratio, not an angle \(tan^{-1}(0^o)\) has no meaning since inverse function take a ratio and give back an angle

OpenStudy (anonymous):

Makes more sense now :) Thanks!